OpenStudy (anonymous):

Solving Systems of Equations by Substitution. -4x + y = 6 -5y - y =21

OpenStudy (anonymous):

-4x = 6 -y x = 3/2 + y/4 ( u make x the subject of the formula in first equation)

sam (.sam.):

-4x + y = 6 .......multiply -1 -5x - y =21 4x - y = -6 -5x - y =21 ---------------- 9x =-27 x=-3 ===================== -4x+y=6 -4(-3)+y=6 y=-6

OpenStudy (anonymous):

is ur qn written correctly